# QAEncoder: Towards Aligned Representation Learning in Question Answering System

### v1

@@ A.7. Missing Figures and Tables
Model Param FIGNEWS(English) FIGNEWS(Arabic) CRUD-RAG(Chinese) FIGNEWS(French) FIGNEWS(Hindi) FIGNEWS(Hebrew)
MRR@8 NDCG@8 MRR@8 NDCG@8 MRR@8 NDCG@8 MRR@8 NDCG@8 MRR@8 NDCG@8 MRR@8 NDCG@8
bge-m3 - 74.3 78.2 77.7 80.5 86.8 88.9 73.4 76.9 58.4 63.8 78.6 81
QAEtxt, $\beta=1.5$ 77.1 80.8 80 82.8 89 90.8 76.7 79.9 62.5 67.5 79.8 82.4
multilingual-e5-small - 70.9 74.7 74 77.1 81.7 84.2 72.8 76.1 66.6 70.1 52.3 57.2
QAEhyb, $\alpha=0.3,\;\beta=0.5$ 73.5 77.5 78.8 81.2 89.1 90.6 77.4 80.1 74.1 77.6 59.4 64
multilingual-e5-base - 74.7 77.9 72.2 75.7 86.1 88.1 70.9 74.3 57.6 62.4 72.6 75.6
QAEemb, $\alpha=0.3$ 77.4 80.7 77.1 80 88.5 90.5 76.6 79.8 61.3 66 77.7 80.4
multilingual-e5-large - 73.8 77.5 76.6 79.9 85.7 88.1 70.4 74 52.7 58.4 73.9 77.1
QAEhyb, $\alpha=0.15,\;\beta=1.25$ 77 80.4 82.1 84.9 89.1 91.1 77.2 80.4 60.3 65 77.6 80.6
gte-multilingual-base - 65.3 69.9 73.2 76.2 82.9 85.6 62.8 67 52.1 57.8 65.9 69.3
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 75.3 79 76.2 79 85.5 88 66.6 70.7 55.8 61.4 73.9 77.2
mcontriever - 32.8 36.3 40.1 44.2 71.7 75.8 34.7 38.6 26.8 31.4 49.2 53.7
QAEhyb, $\alpha=0.45,\;\beta=1.25$ 61.2 65.4 68.2 71.9 88.7 90.5 64.5 68.5 50.3 55.6 69.8 73
bce-embedding-base-v1 - 58.9 63.3 - - 76.8 80.3 - - - - - -
QAEhyb, $\alpha=0.3,\;\beta=0.5$ 66.6 70.6 - - 85.9 88.1 - - - - - -
text2vec-base-multilingual - 38.5 43 27.7 31.4 12.1 13.7 33.2 37.5 15.3 18.9 12.3 14.6
QAEemb, $\alpha=0.75$ 55.2 59.3 51.3 54.8 55.3 58.6 49.1 53.8 35.9 40.6 46.7 51
Table 7. Complete performance comparison of QAEncoder variants on latest datasets FIGNEWS and CRUD-RAG. Higher is better, the best one is bolded. Hyperparameters are optimized simultaneously across the six latest datasets. $n$ indicates the number of predicted queries in QAnaive.
Model Param FIGNEWS(English) FIGNEWS(Arabic) CRUD-RAG(Chinese) FIGNEWS(French) FIGNEWS(Hindi) FIGNEWS(Hebrew)
MRR@8 NDCG@8 MRR@8 NDCG@8 MRR@8 NDCG@8 MRR@8 NDCG@8 MRR@8 NDCG@8 MRR@8 NDCG@8
bge-m3 QAEemb, $\alpha=0.3$ 76.2 80.0 80.0 82.6 88.7 90.6 75.0 78.5 60.9 65.9 78.9 81.7
QAEtxt, $\beta=1.5$ 77.1 80.8 80.1 82.8 89.0 90.8 76.7 79.9 62.5 67.5 79.8 82.4
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 77.3 80.7 80.6 83.3 89.4 91.2 76.4 79.7 61.5 66.3 79.7 82.4
$n=10$ 76.8 79.5 76.9 79.2 85.9 87.8 71.7 73.8 62.2 65.8 68.0 71.3
multilingual-e5-large QAEemb, $\alpha=0.45$ 77.8 81.1 79.7 82.7 89.7 91.4 77.0 79.9 58.2 63.1 78.0 80.9
QAEtxt, $\beta=1.5$ 75.6 79.0 80.7 83.7 88.3 90.4 76.0 79.3 59.7 64.8 77.0 79.7
QAEhyb, $\alpha=0.15,\;\beta=1.25$ 77.0 80.5 82.1 84.9 89.1 91.1 77.2 80.4 60.3 65.1 77.6 80.6
$n=10$ 76.8 79.4 76.5 79.2 84.9 86.8 70.4 73.1 61.5 65.4 69.3 71.8
@@
Table 8. The table illustrates a comprehensive performance comparison of QAEncoder against training-based and document-centric methods on the latest datasets: FIGNEWS and CRUD-RAG. Higher is better, the best one is bolded. Hyperparameters $\alpha,\;\beta$ are optimized simultaneously across the six latest datasets. $n$ denotes the number of pseudo-documents in HyDE. - indicates default or null values.
@@
Model Param FIGNEWS(English) FIGNEWS(Arabic) CRUD-RAG(Chinese) FIGNEWS(French) FIGNEWS(Hindi) FIGNEWS(Hebrew)
MRR@8 NDCG@8 MRR@8 NDCG@8 MRR@8 NDCG@8 MRR@8 NDCG@8 MRR@8 NDCG@8 MRR@8 NDCG@8
mContriever-msmarco - 65.9 70.0 70.1 73.5 85.0 87.4 66.3 70.0 48.6 53.5 69.3 72.6
QAEhyb, $\alpha=0.3,\;\beta=0.75$ 72.1 76.3 77.2 80.2 88.6 90.6 74.2 77.7 58.7 63.5 76.3 79.4
multilingual-e5-large-instruct - 66.7 70.7 74.9 78.1 79.8 82.5 65.8 70.1 48.0 53.1 69.5 72.8
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 75.5 79.4 80.7 83.5 88.9 90.6 75.7 79.2 58.9 64.1 79.3 82.1
mContriever - 32.8 36.3 40.1 44.2 71.7 75.8 34.7 38.6 26.9 31.4 49.2 53.7
HyDE,$n=8$ 24.9 27.6 35.1 40.4 70.7 74.2 25.6 28.6 11.6 14.3 41.8 47.0
multilingual-e5-large - 73.8 77.5 76.6 79.9 85.7 88.2 70.4 74.0 52.7 58.4 73.9 77.1
HyDE,$n=8$ 63.4 67.7 68.2 73.8 81.5 83.9 58.2 62.6 45.5 51.0 65.3 69.6
Figure 6. Ablation on $\alpha$ and $\beta$ hyperparameters for QAEhyb on FIGNEWS(English) dataset.
Figure 7. Ablation on $\beta$ hyperparameter for QAEtxt on FIGNEWS(English) dataset.

### v3

@@ I.2 Challenges of Document-centric Methods
Model Method FIGNEWS(English) FIGNEWS(Arabic) CRUD-RAG(Chinese) FIGNEWS(French) FIGNEWS(Hindi) FIGNEWS(Hebrew)
MRR@10 NDCG@10 MRR@10 NDCG@10 MRR@10 NDCG@10 MRR@10 NDCG@10 MRR@10 NDCG@10 MRR@10 NDCG@10
bge-m3 - 74.4 78.7 77.8 80.9 47.5 48.6 73.5 77.4 58.6 64.4 78.7 81.4
QAEtxt, $\beta=1.5$ 77.2 81 80.2 83.1 51.4 52.5 76.9 80.3 62.7 67.8 80 82.8
multilingual-e5-small - 71 75.1 74.1 77.4 44.6 46.0 66.8 70.4 52.5 57.8 72.9 76.5
QAEhyb, $\alpha=0.3,\;\beta=0.5$ 74.6 78.5 78.9 81.6 50.6 51.6 74.2 77.9 59.6 64.6 77.5 80.4
multilingual-e5-base - 74.8 78.1 72.3 76 47.0 48.2 71.2 75.1 57.8 62.9 72.6 75.8
QAEemb, $\alpha=0.3$ 77.6 81.3 77.2 80.3 51.2 52.3 76.7 80 61.5 66.5 77.7 80.5
multilingual-e5-large - 73.9 77.8 76.7 80.2 46.9 48.3 70.6 74.5 53 59.2 73.9 77.4
QAEhyb, $\alpha=0.15,\;\beta=1.25$ 77.1 80.6 82.2 85.1 51.5 52.7 77.4 80.9 60.6 65.9 77.7 81
gte-multilingual-base - 65.5 70.4 73.4 76.8 45.3 46.8 63 67.4 52.4 58.6 66 69.6
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 75.5 79.5 76.2 79.1 49.4 51.0 66.9 71.6 56 61.9 74 77.6
mcontriever - 32.9 36.7 40.3 44.7 39.2 41.6 35 39.3 27 31.8 49.5 54.4
QAEhyb, $\alpha=0.45,\;\beta=1.25$ 61.4 65.9 68.3 72.1 51.3 52.4 64.7 69.1 50.6 56.4 70.1 73.8
bce-embedding-base-v1 - 59.1 63.8 - - 42.0 44.0 - - - - - -
QAEhyb, $\alpha=0.3,\;\beta=0.5$ 66.8 71.1 - - 49.7 51.0 - - - - - -
text2vec-base-multilingual - 38.7 43.6 27.8 31.9 9.7 10.6 33.7 38.6 15.6 19.9 12.6 15.5
QAEemb, $\alpha=0.75$ 55.4 59.9 51.5 55.4 32.1 34.1 49.3 54.4 36.2 41.5 47.1 52.1
Ablation
bge-m3 - 74.4 78.7 77.8 80.9 47.5 48.6 73.5 77.4 58.6 64.4 78.7 81.4
QAEemb, $\alpha=0.3$ 76.4 80.5 80.1 82.9 51.3 52.4 75 78.5 61 66 79 82
QAEtxt, $\beta=1.5$ 77.2 81 80.2 83.1 51.4 52.5 76.9 80.3 62.7 67.8 80 82.8
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 77.4 81.1 80.6 83.4 51.7 52.7 76.5 80 61.7 66.8 79.8 82.7
multilingual-e5-small - 71 75.1 74.1 77.4 44.6 46.0 66.8 70.4 52.5 57.8 72.9 76.5
QAEemb, $\alpha=0.45$ 74.7 78.5 77 79.8 50.8 51.9 73.6 77.1 58 63.2 77.1 80.1
QAEtxt, $\beta=1.0$ 73.2 77.2 79.2 81.9 49.1 50.5 70.6 74.7 58.7 63.7 77.1 80.2
QAEhyb, $\alpha=0.3,\;\beta=0.5$ 74.6 78.5 78.9 81.6 50.6 51.6 74.2 77.9 59.6 64.6 77.5 80.4
multilingual-e5-base - 74.8 78.1 72.3 76 47.0 48.2 71.2 75.1 57.8 62.9 72.6 75.8
QAEemb, $\alpha=0.3$ 77.6 81.3 77.2 80.3 51.2 52.3 76.7 80 61.5 66.5 77.7 80.5
QAEtxt, $\beta=0.75$ 74.7 78.8 76.4 79.7 50.6 51.8 72.3 76.4 62.5 67.8 77 79.9
QAEhyb, $\alpha=0.3,\;\beta=0.5$ 76.6 80.4 77.4 80.6 51.5 52.5 75.4 79.1 62.7 67.7 77.2 80.2
multilingual-e5-large - 73.9 77.8 76.7 80.2 46.9 48.3 70.6 74.5 53 59.2 73.9 77.4
QAEemb, $\alpha=0.45$ 77.9 81.4 79.8 83 51.9 52.9 77 80 58.3 63.5 78.1 81.2
QAEtxt, $\beta=1.5$ 75.6 79.2 80.9 84.1 51.0 52.3 76 79.4 60 65.5 77 80.2
QAEhyb, $\alpha=0.15,\;\beta=1.25$ 77.1 80.6 82.2 85.1 51.5 52.7 77.4 80.9 60.6 65.9 77.7 81
gte-multilingual-base - 65.5 70.4 73.4 76.8 45.3 46.8 63 67.4 52.4 58.6 66 69.6
QAEemb, $\alpha=0.45$ 69.1 73.6 76.9 79.6 50.5 51.7 67.1 71.2 53 58.5 72.4 75.8
QAEtxt, $\beta=1.5$ 75.7 79.7 75.9 78.9 48.8 50.5 66.1 70.7 56.8 62.5 72.8 76.5
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 75.5 79.5 76.2 79.1 49.4 51.0 66.9 71.6 56 61.9 74 77.6
mcontriever - 32.9 36.7 40.3 44.7 39.2 41.6 35 39.3 27 31.8 49.5 54.4
QAEemb, $\alpha=0.6$ 58.8 64.2 67.2 71.3 51.0 52.1 62.7 66.9 50.4 55.7 69.7 73.2
QAEtxt, $\beta=1.5$ 49.2 54.2 59.5 63.9 46.2 48.2 54.1 58.9 45.1 50.2 60.8 64.6
QAEhyb, $\alpha=0.45,\;\beta=1.25$ 61.4 65.9 68.3 72.1 51.3 52.4 64.7 69.1 50.6 56.4 70.1 73.8
bce-embedding-base-v1 - 59.1 63.8 - - 42.0 44.0 - - - - - -
QAEemb, $\alpha=0.45$ 66.8 71.3 - - 49.3 50.6 - - - - - -
QAEtxt, $\beta=1.5$ 64.3 68.7 - - 47.6 49.2 - - - - - -
QAEhyb, $\alpha=0.3,\;\beta=0.5$ 66.8 71.1 - - 49.7 51.0 - - - - - -
text2vec-base-multilingual - 38.7 43.6 27.8 31.9 9.7 10.6 33.7 38.6 15.6 19.9 12.6 15.5
QAEemb, $\alpha=0.75$ 55.4 59.9 51.5 55.4 32.1 34.1 49.3 54.4 36.2 41.5 47.1 52.1
QAEtxt, $\beta=1.5$ 46.2 50.8 35.5 39.4 12.3 13.6 43.3 47.7 21.3 25.2 18.8 22.1
QAEhyb, $\alpha=0.75,\;\beta=0.5$ 55.1 59.7 50.4 54.4 33.4 35.1 49.1 53.8 34.5 39.6 47.1 51.5
@@
Table 8: Comprehensive retrieval performance on the latest datasets FIGNEWS and CRUD-RAG (Top-k = 10). Higher is better, with the best one bolded. Hyperparameters including QAEncoder variants and weight terms $\alpha,\;\beta$ are optimized simultaneously for six latest datasets. ‘-’ denotes default or null values.
@@
Model Method FIGNEWS(English) FIGNEWS(Arabic) CRUD-RAG(Chinese) FIGNEWS(French) FIGNEWS(Hindi) FIGNEWS(Hebrew)
MRR@10 NDCG@10 MRR@10 NDCG@10 MRR@10 NDCG@10 MRR@10 NDCG@10 MRR@10 NDCG@10 MRR@10 NDCG@10
bg3-m3 QAEemb, $\alpha=0.3$ 76.4 80.5 80.1 82.9 88.8 90.7 75 78.5 61 66 79 82
QAEtxt, $\beta=1.5$ 77.2 81 80.2 83.1 89 90.9 76.9 80.3 62.7 67.8 80 82.8
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 77.4 81.1 80.6 83.4 89.4 91.2 76.5 80 61.7 66.8 79.8 82.7
QAnaive, n=10 76.9 79.9 77.1 79.7 86 88 71.9 74.4 62.3 66.1 68.1 71.7
multilingual-e5-large QAEemb, $\alpha=0.45$ 77.9 81.4 79.8 83 89.8 91.5 77 80 58.3 63.5 78.1 81.2
QAEtxt, $\beta=1.5$ 75.6 79.2 80.9 84.1 88.3 90.5 76 79.4 60 65.5 77 80.2
QAEhyb, $\alpha=0.15,\;\beta=1.25$ 77.1 80.6 82.2 85.1 89.1 91.2 77.4 80.9 60.6 65.9 77.7 81
QAnaive, n=10 77.5 80.3 76.5 79.4 85.1 87.3 70.5 73.5 61.5 65.6 69.4 72.2
Table 9: Complete performance comparison of QAEncoder variants on latest datasets FIGNEWS and CRUD-RAG (Top-k = 10). Higher is better, with the best one bolded. Hyperparameters are optimized simultaneously across the six latest datasets. $n$ indicates the number of predicted queries in QAnaive.
Model Method FIGNEWS(English) FIGNEWS(Arabic) CRUD-RAG(Chinese) FIGNEWS(French) FIGNEWS(Hindi) FIGNEWS(Hebrew)
MRR@10 NDCG@10 MRR@10 NDCG@10 MRR@10 NDCG@10 MRR@10 NDCG@10 MRR@10 NDCG@10 MRR@10 NDCG@10
mcontriever - 32.9 36.7 40.3 44.7 39.2 41.6 35 39.3 27 31.8 49.5 54.4
QAEhyb, $\alpha=0.45,\;\beta=1.25$ 61.4 65.9 68.3 72.1 51.3 52.4 64.7 69.1 50.6 56.4 70.1 73.8
MS 66.1 70.6 70.2 73.7 46.5 47.8 66.4 70.3 48.7 53.9 69.4 72.9
MS + QAEhyb, $\alpha=0.3,\;\beta=0.75$ 72.3 76.8 77.3 80.5 51.2 52.4 74.4 78 59 64.1 76.5 79.8
GPL 68.3 73.6 72.1 75.18 52.8 55.9 70.8 74.3 52.6 55.9 73.9 77.1
QA-RAG 31.1 34.4 42.3 46.8 43.8 45.8 31.8 35.5 29.1 33.9 49.7 55.0
Query2Doc 25.1 29.0 34.5 38.9 35.7 37.2 26.0 29.1 14.2 17.2 39.9 45.0
HyDE 25 27.9 35.7 41.9 36.7 38.7 25.8 28.9 11.8 14.9 42.1 47.8
multilingual-e5-large - 73.9 77.8 76.7 80.2 46.9 48.3 70.6 74.5 53 59.2 73.9 77.4
QAEhyb, $\alpha=0.15,\;\beta=1.25$ 77.1 80.6 82.2 85.1 51.5 52.7 77.4 80.9 60.6 65.9 77.7 81
INS 67 71.4 75 78.2 43.7 45.2 66 70.9 48.2 53.8 69.6 73.3
INS + QAEhyb, $\alpha=0.15,\;\beta=1.5$ 75.6 79.8 80.8 83.7 51.4 52.4 75.8 79.5 59.3 65 79.4 82.3
GPL 75.2 78.9 79.4 82.3 53.6 56.3 73.7 76.8 57.5 62.8 74.8 78.3
QA-RAG 73.3 76.5 72.8 75.6 45.8 46.5 67.2 70.2 55.5 61.3 77.7 80.4
Query2Doc 63.4 68.2 66.3 72.8 42.0 43.1 60.4 65.1 48.4 53.3 62.1 66.0
HyDE 63.6 68.3 68.3 74.1 42.3 43.6 58.4 63.2 45.6 51.3 65.3 69.6
Table 10: The table illustrates a comprehensive performance comparison of QAEncoder against training-based and document-centric methods on the latest datasets: FIGNEWS and CRUD-RAG (Top-k = 10). Higher is better, with the best one bolded. Hyperparameters $\alpha,\;\beta$ are optimized simultaneously across the six latest datasets. $\dagger$ indicates the training-based methods. MS represents fine-tuning on MSMARCO, i.e. the mcontriever-msmarco model; INS represents instruction-tuning, i.e. the multilingual-e5-large-instruct model(Wang et al., 2023a). $\ddagger$ indicates the document-centric methods.
Model Param FIGNEWS(English) FIGNEWS(Arabic) CRUD-RAG(Chinese)
MRR@10 NDCG@10 MRR@10 NDCG@10 MRR@10 NDCG@10
jina-embeddings-v2-small-en - 65.4 69.9 - - - -
QAEhyb, $\alpha=0.15,\;\beta=1.25$ 72.6 76.7 - - - -
jina-embeddings-v2-base-zh - - - - - 41.0 43.1
QAEhyb, $\alpha=0.75,\;\beta=0.75$ 47.5 48.5
jina-embeddings-v2-base-en - 65.3 69.4 - - - -
QAEhyb, $\alpha=0.3,\;\beta=0.5$ 72.4 76.5 - - - -
gte-base-en-v1.5 - 65.6 70.2 - - - -
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 71.3 75.6 - - - -
contriever - 49.6 54.5 - - - -
QAEhyb, $\alpha=0.45,\;\beta=1.25$ 70.7 74.7 - - - -
bge-large-zh-v1.5 - - - - - 42.0 43.7
QAEhyb, $\alpha=0.45,\;\beta=1.25$ 48.6 49.6
bge-large-zh - - - - - 40.6 42.5
QAEhyb, $\alpha=0.45,\;\beta=1.5$ 48.4 49.4
bge-large-en-v1.5 - 66.4 71 - - - -
QAEhyb, $\alpha=0.15,\;\beta=1.5$ 74.3 78.2 - - - -
bge-large-en - 61.9 66.4 - - - -
QAEhyb, $\alpha=0.3,\;\beta=0.75$ 71.3 75.4 - - - -
bge-base-zh-v1.5 - - - - - 43.4 44.9
QAEhyb, $\alpha=0.6,\;\beta=1.0$ 48.6 49.5
bge-base-zh - - - - - 40.3 41.9
QAEhyb, $\alpha=0.6,\;\beta=1.5$ 48.5 49.5
bge-base-en-v1.5 - 66.4 70.5 - - - -
QAEtxt, $\beta=1.0$ 74.1 77.5 - - - -
bge-base-en - 64.9 68.9 - - - -
QAEhyb, $\alpha=0.3,\;\beta=1.5$ 71.7 75.9 - - - -
bce-embedding-base-v1 - 59.1 63.8 - - 42.0 44.0
QAEhyb, $\alpha=0.3,\;\beta=0.5$ 66.8 71.1 - - 47.0 48.3
retriever-multi-base-256 - 33.8 37.4 - - - -
QAEemb, $\alpha=0.75$ 58.4 62.3 - - - -
quora-distilbert-base - 29.8 33.2 - - - -
QAEemb, $\alpha=0.75$ 52.8 57.1 - - - -
Table 11: Retrieval performance of monolingual and bilingual embedding models on the latest datasets FIGNEWS(English) and CRUD-RAG(Chinese). Higher is better, with the best one bolded. ‘-’ denotes default or null values due to mismatch between the language and model.